OpenStudy (anonymous):

x=5t^2 and x=10t. solve for x and t. or what does x=? and t=?

OpenStudy (pfenn1):

Equations given:\[x=5t^2\]\[x=10t\]You can set the first equation equal to the second and solve for t.\[x=5t^2=10t=x\]\[5t^2-10t=5t(t-2)=0\] What is t? What is x?

OpenStudy (anonymous):

how did you get from the first equation to the second equation (5t^2 - 10t - 5t(t-2))?

OpenStudy (anonymous):

i don't quite understand your steps

OpenStudy (pfenn1):

Sorry. Since each of the given equations are equal to x, I set them each equal to one another.\[5t^2=10t\]Subtract 10t from both sides to get\[5t^2-10t=0\] Then I factored out a 5t from the left side\[5t(t-2)=0\]Solve this and 5t=0 and t-2=0. t=2. Plug that back into one of the original equations, like x=10t =10(2)=20. x=20,t=2